how do i factor this completely?..please help...thanks
4a^2+18a-10

Answer by checkley71(8403)

4A^2+18A-10=0
2A^2+9A-5=0
(2A-1)(X+5)=0
2A-1=0
2A=1
A=1/2 SOLUTION
X+5=0
X=-5 SOLUTION